OpenStudy (anonymous):

how would you know what the integral of a graph is if you don't have a calculator to graph and you're not good at envisioning what the graph (x^2 + 1) ^3 looks like? medals given

ganeshie8 (ganeshie8):

so basically you want to sketch the function after evaluating integral ?

ganeshie8 (ganeshie8):

since the given function is a 6th degree polynomial, integrating it would give u a 7th degree polynomial

OpenStudy (anonymous):

i don't necessarily want to sketch it but I want to know what the integral of it is

ganeshie8 (ganeshie8):

oh then its easy, just expand it using binomial theorem

OpenStudy (anonymous):

and then set it equal to 0? so would the integral be 0 to 1?

ganeshie8 (ganeshie8):

why do u want to set it equal to 0 ? whats the exact question ?

OpenStudy (anonymous):

I'm not sure why I did that haha.. I'm just trying to find the integral because the question says: a solid is generated when the region in the first quadrant enclosed by the graph y= (x^2 +1)^3, the line x=1, the x-axis, and the y-axis is revolved around the x-axis. Its volume is found by evaluating what integral?

OpenStudy (anonymous):

So I know that you would multiply pi times the integral of a to b of the function

ganeshie8 (ganeshie8):

oh, okay then you're on right track ! one sec

ganeshie8 (ganeshie8):

Volume by revolving around x-axis : \[\int \limits_0^1 \pi y^2 dx\]

ganeshie8 (ganeshie8):

plugin the value of "y" and evaluate

ganeshie8 (ganeshie8):

\[\int \limits_0^1 \pi y^2 dx =\int \limits_0^1 \pi ((x^2 +1)^3)^2 dx = \pi \int \limits_0^1 (x^2 +1)^6 dx \]

ganeshie8 (ganeshie8):

expand and evaluate

OpenStudy (anonymous):

how did you know that the integral of a to b was 0 to 1 though?

ganeshie8 (ganeshie8):

read the question : ``` when the region in the first quadrant enclosed by the graph y= (x^2 +1)^3, the line x=1, the x-axis, and the y-axis is revolved around the x-axis. ```

ganeshie8 (ganeshie8):

equation of y axis is : x = 0 and you're given : x = 1

ganeshie8 (ganeshie8):